a string of pearls such that 1/3 is lost and of that 1/4th is missing, remaining is 20 then actual number of pearls?
ans. 40

Grade:11

4 Answers

Pankaj Ghildiyal
12 years ago

answer is 48 not 40......... 16 are lost and 12 are missing.....

Sudheesh Singanamalla
12 years ago

Let total Number of pearls be x

number of pearls lost = x/3

number of pearls missing = x/4

remaining = 20

==>  x = x/3 + x/4 + 20

==> x = (4x+3x)/12 + 20

==> x = (7x + 240) / 12

==> 12x = 7x + 240

==> 5x = 240

==> x = 240/5 = 48 !!

 

therefore total number of pearls originally is 48 !

jyoti jadhav
11 years ago
Aneep Agrawal
5 years ago
I think answer is 40 
As it is given that 1/3 is lost and  so remaining perl in string is 2/3 and of that ¼ is missing so
we have           
x/3-2x/3*1/4=20
now we have 
x/3-2x/12=20,
x=40